OpenStudy (anonymous):

Write the explicit formula that represents the geometric sequence -2, 8, -32, 128

OpenStudy (mathmate):

Can you find the common ratio, which is defined as the ratio \(\Large r=\frac{a_{n+1}}{a_n}\)

OpenStudy (anonymous):

I don't think so. How would I do so?

OpenStudy (anonymous):

-4?

OpenStudy (anonymous):

divide 2nd term by first, 3rd term by 2nd etc to find common ratio

OpenStudy (anonymous):

Both equal -4

OpenStudy (anonymous):

that is what @mathmate formula states

OpenStudy (anonymous):

oh okay. I couldn't understand.

OpenStudy (anonymous):

now you do!, so now you have the common ratio: @mathmate will show you, next step

OpenStudy (anonymous):

OKie Dokie.

OpenStudy (anonymous):

Theoretically he will... lol

OpenStudy (anonymous):

Well there is one person who know @acxbox22

OpenStudy (acxbox22):

@BPDlkeme234 wow you are being nice good job i knew you could do it

OpenStudy (anonymous):

Thank you for simplifying thing for me @BPDlkeme234 !

OpenStudy (anonymous):

*things

OpenStudy (anonymous):

Just relying on your expertise here! @acxbox22

OpenStudy (acxbox22):

http://www.regentsprep.org/regents/math/algtrig/atp2/geoseq.htm to find the common ration divide the second term by the first 8/-2=-4 like stated above

OpenStudy (anonymous):

its a simple geometric sequence!

OpenStudy (acxbox22):

r=-4 (r is the common ratio)

OpenStudy (anonymous):

yes, we covered that

OpenStudy (acxbox22):

never hurts to restate it anyway to make it clear @BPDlkeme234

OpenStudy (anonymous):

true!

OpenStudy (anonymous):

So is that the answer then? @acxbox22 @BPDlkeme234

OpenStudy (acxbox22):

yes

OpenStudy (anonymous):

wait for it @Jonnychewy

OpenStudy (anonymous):

no we are not at the formula yet, @acxbox22 is coming up with it..give him some time!

OpenStudy (anonymous):

in the meantime the formula for a geometric sequence is Tn = ar ^n-1

OpenStudy (acxbox22):

xn=ar^(n-1) is the formula where a is the first term and r is the ratio

OpenStudy (acxbox22):

we have those values so we can sub them into the formula to get the answer

OpenStudy (anonymous):

@acxbox22 has already calculated r you see!

OpenStudy (acxbox22):